A sample of argon gas that occupies a volume of 12
Answer #1

argon

V = 12.5 L

T = 0 °C = 273 K

P = 1 atm

contains X moles?

Apply ideal gas law

PV = nRT

n = PV/(RT)

n = (1*12.5)/(0.082*273) = 0.55838 mol of Argon

n = 0.55838 mol of Ar